If 8 years are added to a man's present age and this value is doubled, the result if 100. Find the man's present age
Answer by Cromlix(4381)

Hi there,
Make the man's age = x
(x + 8) * 2 = 100 [* means times]
Now transpose the formula to leave x on its own
x + 8 = 100/2
x + 8 = 50
x = 50 - 8
x = 42
The man's present age is 42 years old.
Hope this helps :-)